A certain kind of sheet metal has, on average, 7 defects per 18 square feet.
Assuming a Poisson distribution, find the probability that a 24 square foot metal sheet has at least 11 defects. Round your answer to three decimal places.

Answer by Boreal(15235)

7/18=x/24
18x=168
x=9.33 (28/3) exact
probability is 0.3342 or 0.334